Find 3 consecutive positive odd integers such that the sun of the squares of the first and second integers is equal to the square of the third integer minus 7
Answer by CubeyThePenguin(3113)

3 positive odd integers: (x-2), x, (x+2)

(x-2)^2 + x^2 = (x+2)^2 - 7
(x^2 - 4x + 4) + x^2 = (x^2 + 4x + 4) - 7
2x^2 - 4x + 4 = x^2 + 4x - 3
x^2 - 8x + 7 = 0
(x - 1)(x - 7) = 0

The three integers have to all be positive, so x = 7 and the three positive integers are 5, 7, and 9.